Main page‎ > ‎Observing Runs‎ > ‎Night Reports‎ > ‎

Night Report 2015-02-25

Observers : Laurent Le Guillou, Sylvain Baumont, Ayan Mitra
Support team : Nicolas Chotard, Pierre François Leget
TAT : K. Maguire, J. Polshaw

We used only EFOSC2, mainly for follow-up observations.

Clear sky at sunset, wind raised from 5m/s to 10m/s, humidity dropped fom 60% to 40%.
Moon passed 1st quarter, 56% illumination, setting at 4:20.
The night was probably not photometric (low atmospheric pressure and some cirrus
towards south-western horizon, downwind, after 4:00).

(times are in UT)

CALIBRATIONS :
 - Evening : EFOSC_bias, EFOSC_FlatSpec, EFOSC_SpecArc
   - 23:40 : EFOSC_SkyFlat gr 
           (tracking issue => only 4 images out of 5, and no time left for i-band)
 - Morning
   - 10:08 : EFOSC_SkyFlat i

SCIENCE OBSERVATIONS :
 --- EFOSC ---
 
 - 00:15  image analysis and focus
 - 00:27  STD L745a, Gr11 Gr13 Gr16, slit 1", airmass 1.1

 - 00:47  OGLE-2015-SN-035 Follow-up, Gr13:2x2700s, slit 1"
        Airmass 1.3, unstable seeing ranging from 1.3" to 2", 86º from the moon.
 - 02:25  PSNJ10544216-2104138 Follow-up, Gr16:2x1800, slit 1.5"
        Airmass 1.2, seeing 1" to 1.5", 99º from the moon.

 - 03:43  SNHiTS15L Classification, Gr13:1800s, slit 1"
        Airmass 1.1, seeing 1" to 1.4", 83º from the moon
        => SN-Ia around 1 week pre-max, z ~ 0.15

 - 04:20  STD LTT3864, Gr11 Gr13 Gr16, slit 1", airmass 1.01
 - 04:31  Photometric Standard Field PG1047, UBVRigrz
        39% humidity, temperature 14º, airmass 1.16

 - 04:43  LSQ14bdq Follow-up, griz:6x200s
        Airmass 1.1, seeing 1" to 1.3"

 - 06:26  Photometric Standard Field PG1323, UBVRigrz
        39% humidity, temperature 14º, airmass 1.12
 - 06:35  Photometric Standard Field PG1047, UBVRigrz
        40% humidity, temperature 14º, airmass 1.23

 - 06:46  PS15ae Follow-up, Gr11:1800s and Gr16:1800s
        Airmass 1.25, seeing 1.1" to 1.4"
 - 08:02  SN2015D Follow-up, Gr11:1500s and Gr16:1500s
        Airmass 1.4, seeing 1.1" to 1.9"

 - 09:12  TCPJ13535979-0755079 Classification attempt, Gr13:1800s, slit 1.5"
        70% humidity, more cirrus, and the telescope cooled down and got out of focus. 
        Airmass 1.1, seeing 1.9" to 1", temperature 9.5º
        => SN-Ic? at z ~ 0.04, but host Halpha gives z ~ 0.009

 - 09:50 STD EG274, Gr11 Gr13 Gr16, slit 1" and 1.5", airmass 1.03


CONCLUSIONS :

  Classification
  * SNHiTS15L            : SN-Ia around 1 week pre-max, z ~ 0.15
  * TCPJ13535979-0755079 : SN-?, z_gal = 0.009 

  Follow-up
  * OGLE-2015-SN-035     Gr13:2x2700s (looks like a IIb extincted by Hydrogen)
  * PSNJ10544216-2104138 Gr16:2x1800s
  * LSQ14bdq             griz:6x200s
  * PS15ae               Gr11:1800s, Gr16:1800s
  * SN2015D              Gr11:1500s, Gr16:1500s

Comments